# The stable trees revisited

### v1

@@ 2.2 Properties of $\widetilde{\sigma}_{t}$
By differentiating the Laplace transform identity in 2.4 at $\lambda=0$ (noting that everything decays rapidly enough for dominated convergence to be valid), we see
@@
Recall that $G(iu)=-\Psi_{L}(u)$. Differentiating the expression for $G$ (again valid because of suitable decay), we see
@@
and by dominated convergence the integral converges to $C_{\alpha}\int_{0}^{\infty}(1-e^{-y})C_{\alpha}y^{-\alpha}\,dy$. Integration by parts reveals that this is equal to $C_{\alpha}\frac{\Gamma(2-\alpha)}{\alpha-1}=\alpha$. ∎
@@
We expect that in fact equality holds for the above, but we only prove an upper bound as it is sufficient for our purposes. We will make use of the following lemma:
@@
$p(-x)$ $=\mathbb{E}\left[\exp\left(\int_{0}^{t}\left(x+\sigma_{s}\right)\,ds\right)p(-\sigma_{s}-x)\right]$
$=\exp(tx)\mathbb{E}\left[\exp\left(\int_{0}^{t}\sigma_{s}\,ds\right)p(-\sigma_{s}-x)\right]$
@@
For $r\geq 0$, define the functions $f_{r}(x,y)=x^{2}\exp(rx)\frac{p(-y-x)}{p(-x)}$. Then
@@
For $r\geq 0$, the (random) function $x\mapsto f_{r}(x,\widetilde{\sigma}_{t})$ is twice continuously differentiable and such that $f_{r}(\cdot,\widetilde{\sigma}_{t})$ along with its first and second partial derivatives in $x$ are vanishing as $x\to\infty$. By Theorem 31.5 of [satobook], $f_{r}(\cdot,\widetilde{\sigma}_{t})$ is in the domain of the generator of $\sigma$, which is the operator $\mathcal{L}$ given by
$\mathcal{L}f(x)=C_{\alpha}\int_{0}^{\infty}f(y-x)y^{-\alpha}dy.$
@@
$\lim_{h\downarrow 0}\frac{1}{h}\mathbb{E}\left[f_{r}(\sigma_{h},\widetilde{\sigma}_{t})\mid\widetilde{\sigma}_{t}\right]=C_{\alpha}\int_{0}^{\infty}x^{2-x}\exp(rx)\frac{p(-\widetilde{\sigma}_{t}-x)}{p(-\widetilde{\sigma}_{t})}dx.$
@@
$\liminf_{h\downarrow 0}\frac{1}{h}\mathbb{E}\left[(\widetilde{\sigma}_{t+h}-\widetilde{\sigma_{t}})^{2}\mid\widetilde{\sigma}_{t}\right]\geq C_{\alpha}\int_{0}^{\infty}x^{2-x}\frac{p(-\widetilde{\sigma}_{t}-x)}{p(-\widetilde{\sigma}_{t})}dx$
@@
$\limsup_{h\downarrow 0}\frac{1}{h}\mathbb{E}\left[(\widetilde{\sigma}_{t+h}-\widetilde{\sigma_{t}})^{2}|\widetilde{\sigma}_{t}\right]$ $\leq C_{\alpha}\inf_{\epsilon>0}\int_{0}^{\infty}x^{2-x}\exp(\epsilon x)\frac{p(-\widetilde{\sigma}_{t}-x)}{p(-\widetilde{\sigma}_{t})}dx$
$=C_{\alpha}\int_{0}^{\infty}x^{2-x}\frac{p(-\widetilde{\sigma}_{t}-x)}{p(-\widetilde{\sigma}_{t})}dx,$
@@
$p(-x)\leq Rx^{\beta}\exp(-C_{2}x^{\frac{\alpha}{\alpha-1}}),\quad p(-x)\geq\frac{1}{R}x^{\beta}\exp(-C_{2}x^{\frac{\alpha}{\alpha-1}})\quad\text{for all $x\geq 0$},$
@@
$\frac{p(-x-y)}{p(-x)}\leq R^{2}\quad\text{for all $x,y\geq 0$}.$ (6)
@@
This bound also depends only on $h$, is finite for each fixed $h$ and is continuous for $h\in(0,1)$. As $h\to 0$, it converges to a constant (one can calculate this constant by integrating the density in (14.37) of [satobook]), and so there exists some $M_{2}$ such that the bound is at most $M_{2}$ for all $h\in(0,1)$. ∎
@@ 2.3 Compactness of the $\mathbb{R}$-tree
Note that by 2.5 we have $\int_{R}^{\infty}\frac{1}{t\mathbb{E}\left[\widetilde{\sigma}_{t}\right]}dt<\infty$ for all large enough $R$. This condition matches that in the compactness conjecture from [abrcompactness], and our proof will follow a similar structure to the proof given there, although we will also require the variance estimate and a few more auxiliary lemmas.
The first lemma is a variation of Lemma 4.7 in [abrcompactness].
@@
Recall the notation for cut points $Y_{i}$ and attachment points $Z_{i}$. We produce a sequence of points $(z_{i},y_{i})_{i\geq 0}\in[0,\infty)^{2}$ as follows. First, $z_{0}=z$. Having constructed $z_{i}$, we define an index $k_{i}=\max\{k:Y_{k}<z_{i}\}$, then set $y_{i}=Y_{k_{i}}$ and $z_{i+1}=Z_{k_{i}}$. That is, $y_{i}$ is the start of the branch containing $z_{i}$, and $z_{i+1}$ is the attachment point for this branch (i.e. $y_{i}^{+}=z_{i+1}$). The path $z_{0}\to z_{1}\to\dots$ gives the geodesic from $y$ back to the root, and terminating upon hitting $\mathcal{T}_{x}$ gives the geodesic from $y$ to this subtree.
@@
First of all, note that $(z_{i},y_{i})$ forms a Markov chain: indeed given values for $(z_{i},y_{i})$ (and $\widetilde{\sigma}$), the law of $(z_{i+1},y_{i+1})$ can be described as follows. First $z_{i+1}$ is sampled according to the normalisation of $d\sigma|_{[0,y_{i})}$. Then $y_{i+1}$ is the rightmost atom to the left of $z_{i}$ in the Poisson process of cuts. Running this procedure as $i$ increases only reveals atoms of the Poisson process from right to left, and so $y_{i+1}$ is conditionally independent of $(z_{0},y_{0}),\dots,(z_{i-1},y_{i-1})$ given $(z_{i},y_{i})$.
@@ 2.4 The mass measure
To prove this lemma, we require the following result about time-dependent Pólya urns (part of which is Theorem 1 of Pemantle [Pemantle]). We defer the straightforward proof to the Appendix.
@@
Take $n=n_{0}$ and pick $\widetilde{S}\subset\mathcal{T}(n_{0}).$ Let $\nu$ be the unique measure on $\mathbb{R}^{+}$ such that $\nu([0,x])=\widetilde{\sigma}_{x}$ for all $x\geq 0$, and note that we can view this as a measure on $\mathcal{T}$ such that $\nu(\mathcal{T}(n))=\widetilde{\sigma}_{Y_{n}}<\infty$ for each $n$. Now set
@@
For each $n$, the subspace $\mathcal{T}_{Y_{n}}$ is compact, hence totally bounded, so we may partition it into finitely many subsets each with diameter at most $1/n$. Write $I^{n}_{1},I^{n}_{2},\dots,I^{n}_{N}$ for this partition, where $N=N(n)$ is the number of pieces. Without loss of generality we may take these partitions to be nested, in the following sense: for any $n<k$, the induced partition of $\mathcal{T}(k)$ from that of $\mathcal{T}(n)$ with pieces $(\pi_{n}^{-1}(I^{n}_{i})\cap\mathcal{T}(k))_{i\leq N(n)}$ is coarser than the partition $(I^{k}_{i})_{i\leq N(k)}$. That is, each piece $I^{k}_{i}$ is a subset of some $\pi_{n}^{-1}(I^{n}_{i})\cap\mathcal{T}(k)$. Now pick arbitrary elements $x^{n}_{i}\in I^{n}_{i}$ for all $n$ and $i$.
@@
has probability 1. We show that the conclusion of the theorem holds on this event. Since $\mathcal{T}$ is compact, the Hausdorff distance $d_{\mathrm{H}}(\mathcal{T},\mathcal{T}(n))$ between $\mathcal{T}$ and $\mathcal{T}(n)$ is finite and tends to zero as $n\to\infty$. In particular, we can bound
@@
exists. We will do this by showing the sequence is Cauchy. Fix any $\varepsilon>0$. Then note that by compactness of $\mathcal{T}$, $g$ is uniformly continuous, so there exists $\delta>0$ such that $d(u,v)\leq\delta$ implies $|g(u)-g(v)|\leq\varepsilon$. Now there exists $n_{0}\in\mathbb{N}$ such that $\delta_{n}<\delta$ for all $n\geq n_{0}$. For any $m>n>n_{0}$, we partition $[N(m)]$ into $N(n)$ pieces given by
@@
Note that the convergence in Theorem 2.14 occurs for all bounded continuous functions simultaneously — a much stronger statement than only having almost sure convergence for each individual bounded continuous function. On this probability 1 event, we can define a map $\mu$ from bounded continuous functions on $\mathcal{T}$ to the reals given by
$\mu(g)=\lim_{n\to\infty}\mu_{n}(g).$
@@
almost surely as $k\to\infty$, for the Gromov–Hausdorff–Prokhorov distance.
@@ 3.1 Reverse Prüfer codes
We now present a bijection between $\mathbb{T}_{n}$ and $n]^{n-1}$, the space of sequences of length n1n-1 taking values in $\{1,2,\dots,n\}$. This is a variant of the classical Prüfer code, and will be useful for constructing random discrete trees. This variant has been studied in [seoshin] and is closely related to the bijection presented in [surveypaper].
@@
and consider the path from $\mathcal{U}_{k-1}$ to $L_{k}$: suppose its vertices are $w_{C_{k-1}},w_{C_{k-1}+1},\dots,w_{C_{k}-1},L_{k}$ in this order (so $w_{C_{k-1}}\in\mathcal{U}_{k-1}$). We repeat this until $\mathcal{U}_{k}=[n]$, i.e. we have exhausted the entire tree. Note that $\mathcal{U}_{k}=C_{k}$ for each $k\geq 1$, and so if we stop after exploring $B$ branches, then $C_{B}=n$. We take as codeword the sequence $(w_{i})_{1\leq i\leq C_{B}-1}$, which has length n1n-1.
@@
We observe that revealing entries in the codeword one at a time gives a way to grow a tree. This is a deterministic construction, but gives a nice way to grow uniformly random trees with a given degree sequence using the fact that the entries of a uniform element of $W_{n,\bm{d}}$ (or indeed $n]^{n-1}$, corresponding to uniformly random trees with no restrictions on individual degrees) have a relatively simple law. Here is an explicit description of this growth process: throughout, we keep track of an ‘active vertex’ where we expand the tree. In the first step, we reveal the root (whose label is $w_{1}$) and mark it as the active vertex. At step $k\geq 2$, we reveal the $k$th entry $w_{k}$ in the codeword, and
@@ 3.2 Relation to Bienaymé trees
We prefer to work with ordered trees, so we use the following random procedure to convert our labelled tree into an ordered tree. First, we map the root $\rho$ to $\varnothing$. Now for any vertex $i$ of $T^{n}$ that is not a leaf, we identify its children and give them a uniformly random ordering, independently of everything constructed so far. Finally, we forget our original labelling. We claim that this gives a conditioned Bienaymé tree with offspring distribution $\xi$.
@@
Fix a rooted ordered tree $t$ of size $n$; we study its probability of arising from our construction. It will be helpful to consider the rooted labelled ordered tree $\widetilde{t}$ that arises in the process just before forgetting the labelling: there are $n!$ possibilities for $\tilde{t}$ that give rise to the final tree $t$, so we study the probability that $\tilde{t}$ arises instead.
@@
Propositions 3.1 and 3.2 show that, provided we have a way to sample $\hat{\bm{D}}$ directly, we can run the following algorithm to generate an unlabelled, unordered conditioned Bienaymé tree. (We note that this procedure is also used in [snakes] to study Bienaymé trees in the case that $\xi$ has finite variance.) Once again we keep track of an active vertex, which we will call $V_{k}$. We will also draw half-edges connected to some of our vertices. Let $H_{k}$ be the set of half-edges after the $k$th step. Most of the interesting behaviour of our algorithm is captured by $V_{k}$ and $H_{k}$, but for technical reasons we must also consider a set of ‘dormant’ vertices, denoted $Z_{k}$. These correspond to vertices created at the ends of branches, whose labels do not show up in the codeword. We will arrange that the active vertex is never dormant.
We begin with the root as our only vertex, which is marked as active, and with no half-edges or dormant vertices. That is, $V_{0}=\rho$, $H_{0}=\varnothing$ and $Z_{0}=\varnothing$. For $k=1,2,\dots$ the $k$th step consists of the following:
@@
  • (Growth event) If $X_{k}=V_{k-1}$, we sample the next entry of $\hat{\bm{D}}$, say $\hat{D}_{i}$. We will show that this is always at least 1. Assuming this for now, we add $\hat{D}_{i}-1$ half-edges to $V_{k-1}$, then add one full edge by creating a new vertex $v$. The dormant set is unchanged from step k1k-1;
  • (Activation event) If $X_{k}=u\in Z_{k-1}$ is a dormant vertex, we sample the next entry of $\hat{\bm{D}}$, say $\hat{D}_{i}$. Once again, we assume for now that $\hat{D}_{i}\geq 1$. We add $\hat{D}_{i}-1$ half-edges to $u$, then add one full edge by creating a new vertex $v$. The previously active vertex $V_{k-1}$ becomes dormant, $u$ ceases to be dormant and the new vertex $v$ becomes the active vertex.
To motivate the choices for the law of $X_{k}$, observe that the process $Z_{k}$ increases by one in every branching event and stays constant in growth and activation events. Thus $Z_{k}$ is equal to the number of repeats among the first $k$ entries of the codeword, and so $n-k+\#Z_{k}$ is the number of unused labels in the codeword. Conditionally given that a branching event does not occur (and the probability of this is $\frac{n-k-\#H_{k-1}}{n-k}$), any given dormant vertex becomes active iff the entry in the underlying codeword matches that of the dormant vertex. This codeword entry is conditionally uniform on the set of unused labels in the codeword so far, and so the conditional probability is $\frac{1}{n-k+\#Z_{k}}$.
@@
$\frac{n-k-\#H_{k-1}}{n-k}=\frac{n-1-\sum_{j=1}^{k-1-\#Z_{k-1}}\hat{D}_{j}}{n-k}\geq 0,$
and since $k-1-\#Z_{k-1}$ is the number of entries of $\hat{\bm{D}}$ revealed so far, equality holds iff we have already revealed all nonzero entries of $\hat{\bm{D}}$. Thus, conditional on the first k1k-1 steps, a growth event at step $k$ has positive probability iff there are still positive entries in $\hat{\bm{D}}$ to reveal.

### v2

@@ 2.2 Properties of $\widetilde{\sigma}_{t}$
We differentiate the Laplace transform identity in 2.4 at $\lambda=0$. By considering the same sequence of approximations used in the proof of 2.2 we see that one can freely exchange derivatives with the Fourier integral, thus obtaining
@@
Recall that $G(iu)=-\Psi_{L}(u)$. By differentiating this (again valid by considering the same approximations) we see
@@
and by dominated convergence the integral converges to $C_{\alpha}\int_{0}^{\infty}(1-e^{-y})y^{-\alpha}\,dy$. Integration by parts reveals that this is equal to $C_{\alpha}\frac{\Gamma(2-\alpha)}{\alpha-1}=\alpha$. ∎
@@
We expect that in fact equality holds for the above, but we only prove an upper bound as it will suffice for our purposes. We will make use of the following lemma:
@@
$p(-x)$ $=\mathbb{E}\left[\exp\left(\int_{0}^{t}\left(x+\sigma_{s}\right)\,ds\right)p(-\sigma_{t}-x)\right]$
$=\exp(tx)\mathbb{E}\left[\exp\left(\int_{0}^{t}\sigma_{s}\,ds\right)p(-\sigma_{t}-x)\right]$
@@
For $r\geq 0$, define the functions $f_{r}(x,y)=x^{2}\exp(rx)\frac{p(-y-x)}{p(-y)}$. Then
@@
For $r\geq 0$, the (random) function $x\mapsto f_{r}(x,\widetilde{\sigma}_{t})$ is twice continuously differentiable and such that $f_{r}(\cdot,\widetilde{\sigma}_{t})$ along with its first and second partial derivatives in $x$ are vanishing as $x\to\infty$. By Theorem 31.5 of [39], $f_{r}(\cdot,\widetilde{\sigma}_{t})$ is in the domain of the generator of $\sigma$, which is the operator $\mathcal{L}$ given by
$\mathcal{L}f(y)=C_{\alpha}\int_{0}^{\infty}\left(f(y+x)-f(x)\right)x^{-\alpha}dx.$
@@
$\lim_{h\downarrow 0}\frac{1}{h}\mathbb{E}\left[f_{r}(\sigma_{h},\widetilde{\sigma}_{t})\mid\widetilde{\sigma}_{t}\right]=C_{\alpha}\int_{0}^{\infty}x^{2-\alpha}\exp(rx)\frac{p(-\widetilde{\sigma}_{t}-x)}{p(-\widetilde{\sigma}_{t})}dx.$
@@
$\liminf_{h\downarrow 0}\frac{1}{h}\mathbb{E}\left[(\widetilde{\sigma}_{t+h}-\widetilde{\sigma_{t}})^{2}\mid\widetilde{\sigma}_{t}\right]\geq C_{\alpha}\int_{0}^{\infty}x^{2-\alpha}\frac{p(-\widetilde{\sigma}_{t}-x)}{p(-\widetilde{\sigma}_{t})}dx$
@@
$\limsup_{h\downarrow 0}\frac{1}{h}\mathbb{E}\left[(\widetilde{\sigma}_{t+h}-\widetilde{\sigma_{t}})^{2}|\widetilde{\sigma}_{t}\right]$ $\leq C_{\alpha}\inf_{\epsilon>0}\int_{0}^{\infty}x^{2-\alpha}\exp(\epsilon x)\frac{p(-\widetilde{\sigma}_{t}-x)}{p(-\widetilde{\sigma}_{t})}dx$
$=C_{\alpha}\int_{0}^{\infty}x^{2-\alpha}\frac{p(-\widetilde{\sigma}_{t}-x)}{p(-\widetilde{\sigma}_{t})}dx,$
@@
$p(-x)\leq R(x^{\beta}\vee 1)\exp(-C_{2}x^{\frac{\alpha}{\alpha-1}}),\quad p(-x)\geq\frac{1}{R}(x^{\beta}\vee 1)\exp(-C_{2}x^{\frac{\alpha}{\alpha-1}})\quad\text{for all $x\geq 0$},$
@@
$\frac{p(-x-y)}{p(-y)}\leq R^{2}\quad\text{for all $x,y\geq 0$}.$ (6)
@@
This bound also depends only on $h$, is finite for each fixed $h$ and is continuous for $h\in(0,1)$. As $h\to 0$, it converges to a constant (one can calculate this constant by integrating the density in (14.37) of [39]), and so there exists some $M_{2}$ such that the bound is at most $M_{2}$ for all $h\in(0,1)$. ∎
@@ 2.3 Compactness of the $\mathbb{R}$-tree
Note that by 2.5 we have $\int_{R}^{\infty}\frac{1}{t\mathbb{E}\left[\widetilde{\sigma}_{t}\right]}dt<\infty$ for all large enough $R$. This condition matches that in the compactness conjecture from [18], and our proof will follow a similar structure to the proof given there, although we will also require the variance estimate and a few more auxiliary lemmas.
The first lemma is a variation of Lemma 4.7 in [18].
@@
Recall the notation for cut points $Y_{i}$ and attachment points $Z_{i}$. We produce a sequence of points $(z_{i},y_{i})_{i\geq 0}\in[0,\infty)^{2}$ as follows. First, $z_{0}=z$. Having constructed $z_{i}$, we define an index $k_{i}=\max\{k:Y_{k}<z_{i}\}$, then set $y_{i}=Y_{k_{i}}$ and $z_{i+1}=Z_{k_{i}}$. That is, $y_{i}$ is the start of the branch containing $z_{i}$, and $z_{i+1}$ is the attachment point for this branch (i.e. $y_{i}^{+}=z_{i+1}$). The path $z_{0}\to z_{1}\to\dots$ gives the geodesic from $z$ back to the root, and terminating upon hitting $\mathcal{T}_{x}$ gives the geodesic from $z$ to this subtree.
@@
First of all, note that $(z_{i},y_{i})$ forms a Markov chain: indeed given values for $(z_{i},y_{i})$ (and $\widetilde{\sigma}$), the law of $(z_{i+1},y_{i+1})$ can be described as follows. First $z_{i+1}$ is sampled according to the normalisation of $d\widetilde{\sigma}|_{[0,y_{i})}$. Then $y_{i+1}$ is the rightmost atom to the left of $z_{i+1}$ in the Poisson process of cuts. Running this procedure as $i$ increases only reveals atoms of the Poisson process from right to left, and so $y_{i+1}$ is conditionally independent of $(z_{0},y_{0}),\dots,(z_{i-1},y_{i-1})$ given $(z_{i},y_{i})$.
@@ 2.4 The mass measure
To prove this lemma, we require the following result about time-dependent Pólya urns (part of which is Theorem 1 of Pemantle [36]). We defer the straightforward proof to the Appendix.
@@
[…]
@@
For each $n$, the subspace $\mathcal{T}_{Y_{n}}$ is compact, hence totally bounded, so we may partition it into finitely many subsets each with diameter at most $1/n$. Write $I^{n}_{1},I^{n}_{2},\dots,I^{n}_{N}$ for this partition, where $N=N(n)$ is the number of pieces. Without loss of generality we may take these partitions to be nested, in the following sense: for any $n<k$, the induced partition of $\mathcal{T}(k)$ from that of $\mathcal{T}(n)$ with pieces $(\pi_{n}^{-1}(I^{n}_{i})\cap\mathcal{T}(k))_{i\leq N(n)}$ is coarser than the partition $(I^{k}_{i})_{i\leq N(k)}$. That is, each piece $I^{k}_{i}$ is a subset of some $\pi_{n}^{-1}(I^{n}_{j})\cap\mathcal{T}(k)$. Now pick arbitrary elements $x^{n}_{i}\in I^{n}_{i}$ for all $n$ and $i$.
@@
has probability 1. We show that the conclusion of the theorem holds on this event. Since $\mathcal{T}$ is compact, the Hausdorff distance $d_{\mathrm{H}}(\mathcal{T},\mathcal{T}(n))$ between $\mathcal{T}$ and $\mathcal{T}(n)$ is finite and tends to zero as $n\to\infty$. It follows that we can bound
@@
exists. We will do this by showing the sequence is Cauchy. Fix any $\varepsilon>0$. By compactness of $\mathcal{T}$, $g$ is uniformly continuous, so there exists $\delta>0$ such that $d(u,v)\leq\delta$ implies $|g(u)-g(v)|\leq\varepsilon$. Now there exists $n_{0}\in\mathbb{N}$ such that $\delta_{n}<\delta$ for all $n\geq n_{0}$. For any $m>n>n_{0}$, we partition $[N(m)]$ into $N(n)$ pieces given by
@@
Note that the convergence in 2.14 occurs for all bounded continuous functions simultaneously — a much stronger statement than only having almost sure convergence for each individual bounded continuous function. On this probability 1 event, we can define a map $\mu$ from bounded continuous functions on $\mathcal{T}$ to the reals given by
$\mu(g):=\lim_{n\to\infty}\mu_{n}(g).$
@@
almost surely, as $k\to\infty$, for the Gromov–Hausdorff–Prokhorov distance.
@@ 3.1 Reverse Prüfer codes
We now present a bijection between $\mathbb{T}_{n}$ and $n]^{n-1}$, the space of sequences of length n1n-1 taking values in $\{1,2,\dots,n\}$. This is a variant of the classical Prüfer code, and will be useful for constructing random discrete trees. This variant has been studied in [40] and is closely related to the bijection presented in [1].
@@
and consider the path from $\mathcal{U}_{k-1}$ to $L_{k}$: suppose its vertices are $w_{C_{k-1}},w_{C_{k-1}+1},\dots,w_{C_{k}-1}$ and $L_{k}$ in this order (so $w_{C_{k-1}}\in\mathcal{U}_{k-1}$). We repeat this until $\mathcal{U}_{k}=[n]$, i.e. we have exhausted the entire tree. Note that $\mathcal{U}_{k}=C_{k}$ for each $k\geq 1$, and so if we stop after exploring $B$ branches, then $C_{B}=n$. We take as codeword the sequence $(w_{i})_{1\leq i\leq C_{B}-1}$, which has length n1n-1.
@@
We observe that revealing entries in the codeword one at a time gives a way to grow a tree. This is a deterministic construction, but gives a nice way to grow uniformly random trees with a given degree sequence (resp. uniformly random trees) using the fact that the entries of a uniform element of $W_{n,\bm{d}}$ (resp. of $n]^{n-1}$) have a relatively simple law. Here is an explicit description of this growth procedure: throughout, we keep track of an ‘active vertex’ from which we expand the tree. In the first step, we reveal the root (whose label is $w_{1}$) and mark it as the active vertex. At step $k\geq 2$, we reveal the $k$th entry $w_{k}$ in the codeword, and
@@ 3.2 Relation to Bienaymé trees
We prefer to work with ordered trees, so we use the following random procedure to convert our labelled tree into an ordered tree. First, we map the root $\rho$ to $\varnothing$. Now for any vertex $i$ of $T^{n}$ that is not a leaf, we identify its children and give them a uniformly random ordering, independently of everything constructed so far. Finally, we forget our original labelling. We claim that this gives a size-conditioned Bienaymé tree with offspring distribution $\xi$.
@@
Fix a rooted ordered tree $t$ of size $n$; we study its probability of arising from our construction. It will be helpful to consider the rooted labelled ordered tree arising in the process just before forgetting the labelling: there are $n!$ possibilities for such a tree that give rise to the final tree $t$, so we fix one such rooted labelled ordered tree $\widetilde{t}$ and study the probability that $\widetilde{t}$ arises instead.
@@
Propositions 3.1 and 3.2 show that, provided we have a way to sample $\hat{\bm{D}}$ directly, we can run the following algorithm to generate an unlabelled, unordered conditioned Bienaymé tree. (We note that this procedure is also used in [2] to study Bienaymé trees in the case that $\xi$ has finite variance.) Once again we keep track of an active vertex, which we will call $V_{k}$. We will also draw half-edges connected to some of our vertices. Let $H_{k}$ be the set of half-edges after the $k$th step. Most of the interesting behaviour of our algorithm is captured by $V_{k}$ and $H_{k}$, but for technical reasons we must also consider a set of ‘dormant’ vertices, denoted $Z_{k}$. These correspond to vertices created at the ends of branches, whose labels do not (immediately) show up in the codeword. We will arrange that no vertex is simultaneously active and dormant.
We begin at time 0 with the root as our only vertex, which is marked as active, and with no half-edges or dormant vertices. That is, $V_{0}=\rho$, $H_{0}=\varnothing$ and $Z_{0}=\varnothing$. For $k=1,2,\dots$ the $k$th step consists of the following:
@@
  • (Growth event) If $X_{k}=V_{k-1}$, we sample the next entry of $\hat{\bm{D}}$, say $\hat{D}_{i}$. We will show that this quantity is always at least 1. Assuming this for now, we create a new vertex $v$, connect it to $V_{k-1}$ by a (full) edge, then add $\hat{D}_{i}-1$ half-edges to $V_{k-1}$. The dormant set is unchanged from step k1k-1;
  • (Activation event) If $X_{k}=u\in Z_{k-1}$ is a dormant vertex, we sample the next entry of $\hat{\bm{D}}$, say $\hat{D}_{i}$. Once again, we assume for now that $\hat{D}_{i}\geq 1$. We create a new vertex $v$, connect it to $u$ by a (full) edge, then add $\hat{D}_{i}-1$ half-edges to $u$. The previously active vertex $V_{k-1}$ becomes dormant, $u$ ceases to be dormant and the new vertex $v$ becomes the active vertex.
To motivate the choices for the law of $X_{k}$, observe that the process $Z_{k}$ increases by one in every branching event and stays constant in growth and activation events. Thus $Z_{k-1}$ is equal to the number of repeats among the first k1k-1 entries of the codeword, and so $n-k+\#Z_{k-1}$ is the number of labels that have not been used in the codeword yet. Conditionally given that a branching event does not occur (and the probability of this is $\frac{n-k-\#H_{k-1}}{n-k}$), any given dormant vertex becomes active iff the entry in the underlying codeword matches that of the dormant vertex. This codeword entry is conditionally uniform on the set of unused labels in the codeword so far, and so the conditional probability is $\frac{1}{n-k+\#Z_{k-1}}$.
@@
$n-k-\#H_{k-1}=n-1-\sum_{j=1}^{k-1-\#Z_{k-1}}\hat{D}_{j}\geq 0,$
and since $k-1-\#Z_{k-1}$ is the number of entries of $\hat{\bm{D}}$ revealed so far, equality holds iff we have already revealed all nonzero entries of $\hat{\bm{D}}$. Thus, conditional on the first k1k-1 steps, a growth event at step $k$ has positive probability iff there are still positive entries in $\hat{\bm{D}}$ to reveal, and an activation event at step $k$ has positive probability iff the same condition holds and there exists at least one dormant vertex.